40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 19:26:35 2013, jamie, Update, Alignment, Adjustment of active TTs and input alignment 
    Reply  Tue Jan 15 20:00:42 2013, Manasa, Update, Alignment, Adjustment of active TTs and input alignment 
Message ID: 7901     Entry time: Tue Jan 15 19:26:35 2013     Reply to this: 7902
Author: jamie 
Type: Update 
Category: Alignment 
Subject: Adjustment of active TTs and input alignment 

[Jamie, Manasa, Jenne]

We started by verifying that the tip-tilts were getting the correct signals at the correct coils, and were hanging properly without touching.

We started with TT2.  It was not hanging freely.  One of the coils was in much further than the others, and the mirror frame was basically sitting on the back side yaw dampers.  I backed out the coil to match the others, and backed off all of the dampers, both in back and the corner dampers on the front.

Once the mirror was freely suspended, we borrowed the BS oplev to verify that the mirror was hanging vertically.  I adjusted the adjustment screw on the bottom of the frame to make it level.  Once that was done, we verified our EPICS control.  We finally figured out that some of the coils have polarity flipped relative to the others, which is why we were seeing pitch as yaw and vice-versa.  At that point we were satisfied with how TT2 was hanging, and went back to TT1.

Given how hard it is to look at TT1, I just made sure all the dampers were backed out and touched the mirror frame to verify that it was freely swinging.  I leveled TT1 with the lower frame adjustment screw by looking at the spot position on MMT1.  Once it was level, we adjusted the EPICS biases in yaw to get it centered in yaw on MMT1.

I then adjusted the screws on MMT1 to get the beam centered at MMT2, and did the same at MMT2 to get the beam centered vertically at TT2.

I put the target at PRM and the double target at BS.  I loosened TT2 from it's base so that I could push it around a bit.  Once I had it in a reasonable position, with a beam coming out at PR3, I adjusted MMT1 to get the beam centered through the PRM target.  I went back and checked that we were still centered at MMT1.  We then adjusted the pitch and yaw of TT2 to get the transmitted beam through the BS targets as clear as possible.

At this point we stopped and closed up.  Tomorrow first thing AM we'll get our beams at the ETMs, try to finalize the input alignment, and see if we can do some in-air locking.

The plan is still to close up at the end of the week.

ELOG V3.1.3-